A two-digit number is 4 times the sum of its digits. The tens digit is 3 less than the units digit. What is the number?
Found 2 solutions by lwsshak3, ewatrrr:
Answer by lwsshak3(11628)

A two-digit number is 4 times the sum of its digits. The tens digit is 3 less than the units digit. What is the number?
**
let u=units digit
t=tens digit
..
t=u-3
10t+u=4(t+u)
10t+u=4t+4u
6t=3u
6(u-3)=3u
6u-18=3u
3u=18
u=6
t=6-3=3
number: 36

Answer by ewatrrr(24785)

 
Hi,
A two-digit number is 4 times the sum of its digits. The tens digit is 3 less than the units digit.
Let x and (x-3) represent the units digit and tens digit respectively
Question states***
10(x-3) + x = 4[(x-3) + x]
Solving for x
11x - 30 = 8x - 12
3x = 18
x = 6, the units digit, the tens digit (x-3) is 3
the number is 36
CHECKING our Answer***
36 = 4*9 = 36